# StateFlow: Enhancing LLM Task-Solving through State-Driven Workflows

### v1

@@ 3.3 A Case Study of StateFlow Design on SQL
4 Experiments
4.1 Setup
We experiment with 2 tasks from the InterCode benchmark Yang et al. (2023a): (1) SQL: The InterCode-SQL adapts the Spider dataset for MySQL, containing 1034 task instances. For each task, a MySQL interpreter is set up with all relevant tables within a docker container. (2) Bash: The InterCode-Bash dataset has 200 task instances curated from the NL2Bash dataset that have 4$\geq$ 4≥ 4 utilities and can be executed in an Ubuntu OS environment. We allow a max of 10 rounds of interaction with the environment. We evaluate on OpenAI GPT-3.5-Turbo and GPT-4-Turbo (the 1106 version) and the temperature is set to 0. Our implementation is based on AutoGen Wu et al. (2023a) [^1] We use AutoGen v0.2.17..
Baselines
For both task, we compare StateFlow with two prompting strategies implemented in the benchmark. (1) ReAct Yao et al. (2022): a few-shot prompting method that prompts the model to generate thoughts and actions. (2) Plan & Solve Wang et al. (2023b): A two-step prompting strategy to first ask the model to propose a plan and then execute it. For the SQL task, we also evaluate a refined version of ReAct, ReAct_Refined, which follows the desired workflow observed in Section 3.3, see Table 7 in appendix for details.
@@ 4.1 Setup
StateFlow 60.83 0.6356 5.38 5.01 5.59 4965 220
SF_Agent 63.73 0.6637 5.67 6.82 3.82 3128 281
GPT-4
Plan & Solve 56.19 0.5793 5.39 1.79 44.7 3065 416
ReAct 60.16 0.6277 5.26 3.87 146.9 12951 419
ReAct_Refined 57.93 0.6104 5.01 2.49 141.06 12377 421
StateFlow 70.41 0.7329 5.84 1.2 49.18 3907 283
SF_Agent 69.34 0.7223 6.11 1.89 36.01 2700 261
@@
StateFlow 37.0 0.8011 3.04 9.95 0.79 3658 148
SF_Agent 36.0 0.8033 3.90 8.74 0.63 2667 232
GPT-4
ReAct 31.5 0.7724 3.86 9.90 20.40 9027 392
@@
StateFlow 37.5 0.8015 2.86 7.27 7.04 3113 135
SF_Agent 39.0 0.8059 2.95 7.85 5.02 1835 225
Metrics
We use 4 metrics provided by the benchmark and we also recorded the LLM usage of each method. (1) Success Rate (SR): A task is considered a success only if the reward is 1. (2) Reward: For SQL, the reward is calculated by Intersection over Union (IoU) of the latest execution output generated against the gold output. For Bash, a customized function is used to evaluate the performance against file system modifications and the latest execution output. (3) Error Rate: percentage of commands that failed. (4) Turns: number of interactions with the environment. (5) Utility: We also recorded the cost, prompt tokens (input), and completion tokens (output) used by each method.
4.2 StateFlow Setup
For both tasks, we follow the ReAct template to prompt the model to generate thought and action at each turn. Each prompt consists of 3 components: instruction, example, and response format. The instruction details what the LLM should perform at the current state, and we use partial example steps from the ReAct examples as demonstrations. The response format asks the LLM to follow the thought-action template to give responses. The prompts are shown in Appendix A.1. SQL: We construct 6 states for the SQL task: Init, Observe, Solve, Error, Verify, End (See Figure 1 and Section 3.3 for a case study). In Init, we always execute the ‘SHOW TABLES’ command. In state Observe, Solve, Verify, Error, when the execution output is an error string, we will transit to state Error. In any of states Solve, Verify, Error, a successful ‘DESC’ will transit to Solve; a successful ‘SELECT’ will transit to Verify. Bash: For the bash task, we define a StateFlow model consists of 5 states Init, Solve, Verify, End, Error. The states are similar to SQL, and the transition only depends on whether the execution is successful or not (Figure 1). We evaluate both StateFlow and SF_Agent with the same setup. In StateFlow, each instruction sent to the model will be preserved in context history. In SF_Agent, the same instruction is used to prompt each agent and is not visible to other agents.
4.3 Results
On both GPT-3.5 and GPT-4, StateFlow and SF_Agent have a dominating advantage both in terms of performance and efficiency over other prompting methods. SQL: Specifically on GPT-3.5, our refined ReAct version increases the success rate by 7% over the original ReAct prompt, but SF_Agent can further improve over the refined version by 6%, with 5$\times$ less cost. Compared to Plan & Solve, SF_Agent uses 1.6$\times$ more cost but improves the SR by 17%. We note that ReAct_Refined and our methods follow similar workflow and they both have low error rates. On GPT-4-Turbo, StateFlow can improve over ReAct by 10% in SR, but has 3$\times$ less cost. Bash: Similarly, StateFlow and SF_Agent outperform other methods while being efficient at interacting with the environment and utilizing LLM computation on both GPT-3.5 and GPT-4. Comparing StateFlow and SF_Agent: Overall, StateFlow and SF_Agent have similar performance in terms of SR and reward. Since StateFlow has a cumulative set of prompts in the history, it has a much higher prompt token usage, resulting in a higher cost. However, with the instructions in history, StateFlow is more efficient in interacting with the environment, with a lower number of turns and error rates.
4.4 Ablation Study
SR Turns Error Cost \$
% $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
StateFlow 60.83 5.38 5.01 5.59
No_Verify 58.60 5.29 4.90 6.65
No_Error 55.31 5.64 11.23 6.50
No_Observe 59.30 5.62 14.65 7.47
SF_Agent 63.73 5.67 6.82 3.82
No_Verify 62.28 5.18 5.96 3.68
No_Error 58.80 5.72 11.58 4.05
No_Observe 57.83 6.00 16.95 4.64
Table 3: Ablation of states on the InterCode SQL dataset with GPT-3.5-Turbo. Best metrics in Bold. Second-best is Underlined. For both StateFlow and SF_Agent, we remove three different states and show the results. Full table in appendix Table 4.
To understand how different states contribute to the accuracy in StateFlow, we perform additional ablations and analysis with the SQL task (See Table 3). (1) We remove the Observe state. Note that in the original prompt for state Solve, we also instruct the model to call ‘DESC’ if necessary, so the model can still explore tables, but not with an explicit state and instruction to perform this action. (2) We remove the error state and rely on the verify state to correct mistakes. (3) We remove the verify state and add a sentence in Solve to prompt the model to submit when finished.
From the table, we can see that the removal of the three different states all result in a drop in performance. When Verify state is removed, both StateFlow and SF_Agent have a lower cost and error rate, matching the idea that more corrections to the results will be performed with an explicit Verify state. Removal of the Error state leads to a drop of 5% in SR for both views, and an increase in turns, error rate, and cost, showing that the Error state plays an important role in the workflow. Removal of the Observe state has minimal impact on the performance of StateFlow, but results in a great drop in performance in SF_Agent. We hypothesize that, with repetitive instructions in context, the LLM can better capture the idea of calling ‘DESC’ to explore tables when getting errors. The trade-off is a higher cost incurred by LLM generation.
5 Related works
Different frameworks have been proposed for LLM inference Sumers et al. (2023); Sel et al. (2023); Ning et al. (2023); Zhang et al. (2023); Wu et al. (2022). Tree of thoughts (ToTs) Yao et al. (2023) models the reasoning process as a tree. While each node in the tree represents a thought, multiple sequential thoughts are generated and evaluated to determine the next node. Tree-of-Thought (ToT) Long (2023) also models the thoughts as trees, but relies on a rule-based verifier to determine if a thought is valid and performs refining or backtracking based on a controller. Graph of Thoughts Besta et al. (2023) models the process as a directed graph and defines 3 types of transformations at a node in the graph: aggregation, refining, and generation. In parallel to our study, State Machine of Thoughts (SMOT) Liu and Shuai (2023) also models thoughts as state machines, but uses one planning agent to manage state transition and one action agent to perform actions. While existing methods model the thought generation process of LLMs as graphs or trees, our framework models a general task-solving process, actively considering the incorporation of feedback from tools and environments. The model of state machines allows a dynamic design to solve more complex tasks, where any state can be connected with the definition of state transitions. We note that the step-wise voting strategy from TOTs Yao et al. (2023) can easily be applied to StateFlow. When we use LLM to general outputs in a state, we can generate several outputs and select the best one.
It has been shown that utilizing LLMs to perform self-evaluation and self-correction can refine intermediate steps during the complex reasoning process Weng et al. (2022); Xie et al. (2023); Paul et al. (2023); Shinn et al. (2023); Kim et al. (2023); Zhou et al. (2023). This strategy is also used in coding Chen et al. (2023a) and computer tasks Kim et al. (2023). In our experiment, we also employ a Verify state that asks the LLM to verify the problem has been solved, and a Error state to correct errors.
LLMs have been used to interact with environments Deng et al. (2023); Yang et al. (2023a) and tools Paranjape et al. (2023); Gou et al. (2023); Schick et al. (2023); Gao et al. (2023); Yang et al. (2023b). Wu et al. (2023b); Gao et al. (2022) improves math solving process through Python execution, and Shi et al. (2023) augments LLMs with a retrieval model for knowledge-intensive tasks. Essentially, the interaction with tools can viewed as a ‘single-use’ of environments Sumers et al. (2023).
LLMs are becoming a promising foundation for developing autonomous agents Xi et al. (2023); Wang et al. (2023a); Wu et al. (2023a); Li et al. (2023); Hong et al. (2023). AutoGen Wu et al. (2023a) offers an open-source platform for developing LLM-based agents. MetaGPT Hong et al. (2023) proposes a multi-agent framework for software development and CAMEL Li et al. (2023) proposes a framework for autonomous agent cooperation. We also propose SF_Agent, an agent view that allows agents to perform different actions at different states.
6 Conclusion
In this paper, we propose StateFlow, a novel problem-solving framework to use LLMs for complex, multi-step tasks with enhanced efficiency and control. By conceptualizing the task-solving process as a state machine, we distinguish between “process grounding” (via state transitions and states) and “sub-task solving” (through actions within a state), enhancing control and interpretability of the task-solving procedure. StateFlow outperforms existing methodologies in terms of success rates and cost efficiency, as evidenced by its application to the SQL and Bash tasks in the InterCode benchmark.
An intriguing avenue for further research lies in the automation of StateFlow model construction, leveraging LLMs to dynamically generate and refine workflows. Further, the idea of employing active learning strategies to iteratively adjust or "train" a StateFlow, adding or removing states automatically based on task performance, presents a promising path toward maximizing efficiency and adaptability in complex task solving.
Limitations
Similar to other frameworks, StateFlow also requires humans to have a good understanding of a given task and build the model and prompts. Future work would be to automate the building process, employ LLMs to observe tasks and construct the model, and use automatic prompting methods to refine the prompts. Also, in comparison to GoT, our current design of StateFlow cannot perform parallel actions and then merge them. This could be done in future work to leverage asynchronous system designs.
Acknowledgements
We would like to thank Tiffany Ling from MathGPTPro for her assistance in creating the figures.
Appendix A Appendix
A.1 Experiment details
See table 5 for the instructions used for SQL StateFlow model and Table 2 for instructions used for Bash. See Figure 5 We include two examples of ReAct and ReAct_Refined in Table 7 for comparison.
Figure 4: StateFlow model for the ablation study.
Figure 5: Example of StateFlow on Bash task.
@@ A.1 Experiment details
StateFlow 60.83 0.6356 5.38 5.01 5.59 4965 220
No_Verify 58.60 0.6169 5.29 4.90 6.65 5981 225
No_Error 55.31 0.5747 5.64 11.23 6.50 5788 253
No_Observe 59.30 0.6197 5.62 14.65 7.47 6696 264
SF_Agent 63.73 0.6637 5.67 6.82 3.82 3128 281

### v2

@@ 4.1 InterCode Benchmark
StateFlow Setup. For both tasks, we prompt the model to generate thought and action at each turn. Each prompt consists of 3 components: (1) instruction: details of what the LLM should perform at the current state; (2) examples: partial thought or action steps from the ReAct examples as demonstrations; (3) response format: explain the thought-action template. These prompts are put in the system message of each LLM agent and are not visible to other agents. SQL: We construct 6 states for the SQL task (See Figure 1 and Section 3.3 for a case study). In Init, we always execute the ‘SHOW TABLES’ command. In state Observe, Solve, Verify, Error, when the execution output is an error string, we will transit to state Error. In any of states Solve, Verify, Error, a successful ‘DESC’ will transit to Solve; a successful ‘SELECT’ will transit to Verify. In Verify, we use LLM to self-evaluate, which is proven useful by Weng et al. (2022); Xie et al. (2023). Bash: For the bash task, we define a StateFlow model consists of 5 states Init, Solve, Verify, End, Error. The states are similar to SQL, and the transition only depends on whether the execution is successful or not (Figure 1). See details in Appendix A.1
GPT-3.5 GPT-4
[…]
Plan & Solve 47.68 4.31 12.5 2.38 56.19 5.39 1.79 44.7
ReAct 50.68 5.58 16.3 17.7 60.16 5.26 3.87 147
ReAct_Refine 57.74 5.47 3.82 18.1 57.93 5.01 2.49 141
StateFlow 63.73 5.67 6.82 3.82 69.34 6.11 1.89 36.0
Results and analysis. On both SQL and Bash tasks, StateFlow outperforms other prompting methods in terms of both performance and efficiency. SQL: On GPT-3.5, our refined ReAct version increases the success rate by 7% over the original ReAct prompt. But StateFlow can further improve over the refined version by 6%, with 5$\times$ less cost. We note that the big difference in cost mainly comes from the prompt token use. The ReAct prompt has 2043 tokens with 4 example trajectories, while the longest instruction from StateFlow has only 400 tokens. The LLM is called iteratively with the whole prompt, thus the difference in cost accumulates. Compared to Plan & Solve, StateFlow uses 1.6$\times$ more cost but improves the SR by 17%. We note that ReAct_Refined and our methods follow similar workflow and they both have low error rates. On GPT-4-Turbo, StateFlow can improve over ReAct by 10% in SR but has 3$\times$ less cost. Bash: On the bash task, StateFlow outperforms other methods while being efficient at interacting with the environment. Switching to GPT-4-Turbo has little effect on the methods, where the two baselines even suffer from a decrement in accuracy. More analysis and results are in Appendix A.2 and A.3.
GPT-3.5 GPT-4
[…]
Plan & Solve 23.5 4.98 25.8 0.74 20.5 5.15 21.0 9.59
ReAct 32.5 5.52 13.2 3.28 31.5 3.86 9.90 20.40
StateFlow 36.0 3.90 8.74 0.63 39.0 2.95 7.85 5.02
SR Turns Error Cost
% $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
StateFlow 63.73 5.67 6.82 3.82
No_Verify 62.28 5.18 5.96 3.68
No_Error 58.80 5.72 11.6 4.05
No_Obsrve 57.83 6.00 17.0 4.64
Figure 4: Ablation of states on the InterCode SQL dataset with GPT-3.5-Turbo. Best metrics in Bold. Second-best is Underlined.
Ablation of States. To understand how different states contribute to the accuracy in StateFlow, we perform additional ablations and analysis with the SQL task (See Table 4). (1) We remove the Observe state. Note that in the original prompt for state Solve, we also instruct the model to call ‘DESC’ if necessary, so the model can still explore tables, but not with an explicit state and instruction to perform this action. (2) We remove the error state and rely on the verify state to correct mistakes. (3) We remove the verify state and add a sentence in Solve to prompt the model to submit when finished. The table shows that removing any of the states results in a drop in performance. When Verify state is removed, StateFlow has the lowest cost and error rate, matching the idea that more corrections to the results will be performed with an explicit Verify state. Removal of the Error state leads to a drop of 5% in SR, and an increase in turns, error rate, and cost, showing that the Error state plays an important role in the workflow. Removal of the Observe results in the lowest SR and highest cost, showing that ‘Observe‘ is the most important state.
4.2 ALFWorld
ALFWorld (Shridhar et al., 2020) is a synthetic text-based game implemented in the TextWorld environments (Côté et al., 2019). It contains 134 tasks across 6 distinct task types: move one or two objects (e.g., put one/two cellphone in sofa), clean/cool/heat an object (e.g., clean/cool/heat some apple and put it in sidetable) and examine an object with lamp (e.g., look at bowl under the desklamp). The agent is required to navigate around a household setting and manipulate objects through text actions (e.g., go to desk 1, take soapbar from toilet 1), and the environment will give textual feedback after each action. We experiment with GPT-3.5-Turbo (1106). For all experiments, we follow AutoGen (Wu et al., 2023a) to use the BLEU metric to map output to the valid action with the highest similarity. Our implementation is also based on AutoGen [^1] We use AutoGen v0.2.17.. More details are in Appendix B.1.
Figure 5: The StateFlow model for ALFWorld. For Plan, we call the LLM directly. For other states (except Init and End), we first call LLM with an instruction and then call the environment to get feedback. In state Pick, when the correct object is picked, we transit to different states based on task type. For states Pick, Process, FindLamp, UseLamp, Put, we stay in the current state if the corresponding task is not completed, represented by gray semi-circle arrows.
Baselines. We evaluate with: 1. ReAct (Yao et al., 2022b): We use the two-shot prompt from the ReAct. Note there is a specific prompt for each type of task. 2. ALFChat (2 agents) (Wu et al., 2023a): A two-agent system setting from AutoGen consisting of an assistant agent and an executor agent. ALFChat is based on ReAct, which modifies the ReAct prompt to follow a conversational manner. 3. ALFChat (3 agents): Based on the 2-agent system, it introduces a grounding agent to provide commonsense facts whenever the assistant outputs the same action three times in a row.
[…]
ReAct 83.3 36.6 53.6 58.7 63 41.2 55.5 6.6
ALFChat (2 agents) 87.5 60.2 44.9 65.1 38.9 43.1 58.2 6.9
ALFChat (3 agents) 84.7 60.2 69.6 77.8 68.5 41.2 67.7 6.1
StateFlow 91.7 83.9 85.5 79.4 92.6 62.7 83.3 2.6
Table 3: Performance and cost of StateFlow and other methods on ALFWorld benchmark with GPT-3.5-Turbo. We report average success rate of 3 attempts.
StateFlow Setup. The StateFlow model is shown in Figure 5. The process starts at Plan, where a plan to solve the task is generated. We note a similar planning is also used in the ReAct prompting. Then, we transit to Pick to instruct the model to search for the target object and take it. We stay in Pick until the target object is picked.
We identify the target object by calling an LLM at the beginning and use it as ground truth. If a wrong object is picked, we go to the Pick_Error state, where the wrong object is put down. If the correct object is picked, we transit to the next state (Process, Put or FindLamp) based on the task type. We follow the ReAct template to prompt the model to generate thought and action each time. Also, we use task-specific planning examples and instructions in Plan and Process. See details in Appendix B.1.
Figure 6: StateFlow and ReAct integrated with Reflexion. StateFlow can further be improved with Reflexion, with much less cost incurred than ReAct.
Results and analysis. Table 3 shows the results for the ALFWorld benchmark. We record the accuracy for each type of task and also the cost for LLM inference. We can see that StateFlow achieves the best performance on all 6 tasks, and significantly outperforms all baseline methods on the whole dataset. It improves over ReAct by 28% and ALFChat (3 agents) by 15%. At the same time, StateFlow uses 2.5x less cost. With StateFlow, we decompose a long prompt into shorter but more concise prompts to be used when entering a state. Thus, we reduce the prompt tokens used while making the model focus on a sub-task for better responses. We refer to Appendix B.2 for more analysis of the results. In Appendix B.3, we show that further decomposition of the states in StateFlow can increase the success rate to 88.8%, with 15% less cost.
Integration with Reflexion. We incorporate StateFlow with Reflexion (Shinn et al., 2023), showing that StateFlow can be combined with iterative refining methods to improve its performance. Reflexion reflects from previous unsuccessful trials and stores the feedback in memory for subsequent trials. We can either use ReAct or StateFlow as the basic executor. We run StateFlow +Reflexion and ReAct+Reflexion for 6 iterations, until both ReAct and StateFlow stop performance improvement. In Figure 6, StateFlow +Reflexion further improves from 84% to 94.8%, with the total cost increased from \$2.9 to \$8.6. While ReAct+Reflexion improves from 55.2% to 74.6%, the total cost for it increased from \$7.1 to \$27.9.
5 Related work
Different prompting frameworks have been proposed to enhance LLM reasoning processes (Zhang et al., 2023b; Wu et al., 2022; Sel et al., 2023; Ning et al., 2023; Zhou et al., 2022; Zhang et al., 2023a). Tree of thoughts (ToTs) (Yao et al., 2023) models the reasoning process as a tree and employs DFS or BFS search to explore sequential thoughts. Tree-of-Thought by (Long, 2023) models the thoughts as trees but relies on a rule-based verifier to determine if a thought is valid and performs refining or backtracking based on a controller. Graph of Thoughts (Besta et al., 2023) models the process as a directed graph and defines 3 types of transformations at a node in the graph: aggregation, refining, and generation. These frameworks have better control over LLM’s intermediate steps, but they are not well-designed to consider LLM workflows with external tools and environments. StateFlow considers external feedback and also allows the design of complex patterns for more difficult tasks, where any state can be connected with the definition of state transitions. The step-wise search from ToTs (Yao et al., 2023) can easily be applied to StateFlow. When we call LLM in a state, we can generate several responses and employ an evaluator to select the best one.
LLMs have been used to interact with environments (Deng et al., 2023; Yao et al., 2022a; Shridhar et al., 2020) and tools (Paranjape et al., 2023; Gou et al., 2023; Schick et al., 2023; Gao et al., 2023; Yang et al., 2023b; Zhang et al., 2024). ReAct (Yao et al., 2022b) uses a few-shot prompting strategy that generates the next action based on past actions, and observations, which have been proven effective. Follow-up works employ iterative refining to improve from previous trials (Sun et al., 2024; Prasad et al., 2023). Reflexion (Shinn et al., 2023) and Self-Refine (Madaan et al., 2023) generate reflections from the past to improve future trials. In parallel to our work, Liu & Shuai (2023) adapts a state machine to record and learn from past trajectories. Extending from ToTs and RAP (Hao et al., 2023)(Zhou et al., 2023) proposes an LLM-based tree search incorporating reflection and feedback from the environment. These methods incur additional costs from the interactions or searching processes. StateFlow is orthogonal to these methods, and some of them can be used on top of StateFlow to further improve performance.
LLMs are becoming a promising foundation for developing autonomous agents (Xi et al., 2023; Wang et al., 2023a; Wu et al., 2023a; Li et al., 2023; Hong et al., 2023; Sumers et al., 2023). AutoGen (Wu et al., 2023a) offers an open-source platform for developing LLM-based agents. MetaGPT (Hong et al., 2023) proposes a multi-agent framework for software development and CAMEL (Li et al., 2023) proposes a framework for autonomous agent cooperation. In StateFlow, specialized agents with different instructions are used in different states.
6 Conclusion
In this paper, we propose StateFlow, a novel problem-solving framework to use LLMs for complex, multi-step tasks with enhanced efficiency and control. StateFlow grounds the progress of task-solving with states and transitions, ensuring clear tracking and management of LLMs’ responses and external feedback. We can define a sequence of actions within each state to solve a sub-task. StateFlow requires humans to have a good understanding of a given task and build the model and prompts. An intriguing avenue for further research lies in the automation of StateFlow model construction and prompting writing, leveraging LLMs to dynamically generate and refine workflows. Further, the idea of employing active learning strategies to iteratively adjust or ”train” a StateFlow, adding or removing states automatically based on task performance, presents a promising path toward maximizing efficiency and adaptability in complex task solving.
Acknowledgements
We would like to thank Hanjun Dai and Eric Zelikman for their reviews and helpful feedback. We also thank Yu Tong Tiffany Ling from MathGPTPro for creating demonstration figures.
Appendix A InterCode
A.1 Experiment Details
For ReAct and Plan & Solve, we use the code from InterCode repository [^2] https://github.com/princeton-nlp/intercode. The StateFlow models for the ablation study on the InterCode SQL task are in Figure 9 and the full metrics for the ablation study are shown in Table 7. See Figure 10 for a bash example with StateFlow. We include also two examples of ReAct and ReAct_Refined in Table 18 and  19 for comparison.
For the InterCode benchmark, we recorded different metrics provided by the benchmark and we also recorded the LLM usage of each method. Additional metrics: (1) Reward: For SQL, the reward is calculated by Intersection over Union (IoU) of the latest execution output generated against the gold output. For Bash, a customized function is used to evaluate the performance against file system modifications and the latest execution output. (2) Token Count: We also recorded the prompt tokens (input), and completion tokens (output) used by each method. See Table 5 and Table 6 for detailed comparisons.
See Table 14 and 15 for the instructions used for SQL StateFlow model and Table 16 and 17 for instructions used for Bash. Each instruction is constructed with a uniform prompt that introduces the overall environment and a specific prompt that instructs the model to act. We follow the InterCode benchmark to use the chat-based models as text completions. All the instructions are put in the system message of the model. The task is entered as the user message, and new responses from the model and environment will be appended to that user message, following a completion manner.
Code is available at https://github.com/kaebvcidn/StateFlow.
A.2 Additional Analysis
The additional analysis below is based on results with the GPT-3.5-Turbo model.
SQL. The SQL dataset consists of different levels of difficulties (See Table 4). The success rate drops with harder tasks, and the difference in SR between easy and extra hard tasks is as great as 50% with StateFlow. Harder queries usually pose several constraints on the data and require looking across tables and joining information from them (Extra hard example: “Which distinctive models are produced by maker with the full name General Motors or weighing more than 3500?”. Easy example: “Give the city and country of the Alton airport.”). However, StateFlow greatly improves over other baselines on hard and extra tasks, leading to 20% improvement compared to ReAct. Since harder tasks require information across tables, they are more likely to result in errors. We collected the states traversed for tasks that are solved successfully, and found that only 9% of the easy tasks go over state Error, while 20% of the extra hard tasks have state Error visited. This indicates that the state Error in StateFlow plays an important role in the performance gain in hard tasks.
Easy Medium Hard Extra All
Plan & Solve 80.6 49.1 29.1 13.2 47.7
ReAct 72.2 57.6 35.6 15.7 50.7
ReAct_Refined 77.0 66.1 44.8 19.9 57.7
StateFlow 87.9 62.9 59.8 36.7 63.7
Table 4: Success Rate of different level of difficulties on InterCode SQL with GPT-3.5-Turbo.
[…]
A.3 Additional Results
SF_Chat  We present the results of SF_Chat, an alternative version of StateFlow, in Table 5 and 6. Instead of creating individual LLM agents with specific instructions, we directly append that instruction to context history, imitating a user’s reply in a conversation. For this alternative, we construct the context history in a conversational manner. The observations and instructions are appended as user messages, and replies from models are appended as assistant messages. We can see that SF_Chat has a similar performance to StateFlow. By directly appending the instructions, SF_Chat takes fewer turns to finish the task and has a lower error rate than StateFlow. In trade-off, the cost of SF_Chat is higher. We note that SF_Chat might not be suitable for tasks that require many turns of interactions (e.g., ALFWorld), because the cost would be high with the instruction prompts accumulated in the context history.
@@ A.3 Additional Results
Try Again 56.38 0.5762 7.62 34.73 6.61 6098 145
StateFlow 63.73 0.6637 5.67 6.82 3.82 3128 281
SF_Chat 60.83 0.6356 5.38 5.01 5.59 4965 220
GPT-4 Plan & Solve 56.19 0.5793 5.39 1.79 44.7 3065 416
ReAct 60.16 0.6277 5.26 3.87 147 12951 419
ReAct_Refined 57.93 0.6104 5.01 2.49 141 12377 421
StateFlow 69.34 0.7223 6.11 1.89 36.0 2700 261
SF_Chat 70.41 0.7329 5.84 1.20 49.2 3907 283
Table 5: Results of the Intercode SQL with GPT-3.5 and GPT-4 with all metrics. We also include SF_Chat, an alternative of StateFlow, and another baseline Try Again with an oracle setting(). Best metrics of each model is in Bold. Second-best is Underlined.
@@
Try Again 49.5 0.8453 6.88 19.54 0.83 3833 159
StateFlow 36.0 0.8033 3.90 8.74 0.63 2667 232
SF_Chat 37.0 0.8011 3.04 9.95 0.79 3658 148
GPT-4 ReAct 31.5 0.7724 3.86 9.90 20.40 9027 392
@@
StateFlow 39.0 0.8059 2.95 7.85 5.02 1835 225
SF_Chat 37.5 0.8015 2.86 7.27 7.04 3113 135
Table 6: Results of the Intercode Bash with GPT-3.5 and GPT-4 with all metrics. We also include SF_Chat, an alternative of StateFlow, and another baseline Try Again with an oracle setting().
Try Again  We also include results of another baseline “Try Again” from InterCode Benchmark with GPT-3.5-Turbo. Try Again is an iterative feedback setup from InterCode to mimic human software development (Yang et al., 2023a). In this setup, the model can receive a ground-truth reward from the environment at each execution of the command and stops when the task is solved correctly or reaches max turns. Then the max reward from all the executions is retrieved. We note that this is an oracle setting not used in StateFlow and other baselines. In our setting, we use the model to determine when to stop and submit the answer, and only the result before submission is evaluated. From Table 5, we can see that Try Again doesn’t work well in the SQL task, and that the performance is slightly worse than our refined version of ReAct. It also has a high error rate of 34.73%. However, with the bash task, Try Again significantly outperforms other methods. This discrepancy indicates the difference in the nature of the two tasks. In the SQL task, the reward is mostly 0 or 1. The observation commands such as ”DESC” and a wrong ”SELECT” command receive 0, and there are only a few cases where the ”SELECT” command is partially correct to receive a partial reward. In this case, the reward signal is not very useful. However, the bash tasks from InterCode are explicitly selected with utilities $\geq$ 4 (require several commands), and each correct command can receive a partial reward. Thus, the reward signal can help the model understand how much progress it has made and provide guidance, leading to a significant improvement in performance.
@@
StateFlow 63.73 0.6637 5.67 6.82 3.82 3128 281